How much of a 60%-acid solution must be mixed with 70 gallons of a 16%-acid solution to obtain a
solution that is 50% acid?

Answer by stanbon(75887)

How much of a 60%-acid solution must be mixed with 70 gallons of a 16%-acid solution to obtain a solution that is 50% acid?
----
Equation:
acid + acid = acid
0.16*70 + 0.60x = 0.50(70+x)
Multiply thru by 100 to get:
16*70 + 60x = 50*70 + 50x
10x = 34*70
x = 34*7
x = 238 gallons (amt of 60% solution needed for the mixture)
=============================================
Cheers,
Stan H.